OpenStudy (anonymous):

what force must be applied to the end of a 2.0 meter long crowbar in order to lift a 500N rock if the fulcrum of the bar is .5meters from the rock?

OpenStudy (anonymous):

use force multiplication or torque concept and the answer is 500/3 N

OpenStudy (kira_yamato):

F = 500x0.5 / 1.5 F = 500/3 = 166 2/3 N